### v2

# Swampland Statistics for Black Holes

Abstract
In this work, we approach certain black hole issues, including remnants, by providing a statistical description based on the weak gravity conjecture in the swampland program. Inspired by the Pauli exclusion principle in the context of the Fermi sphere, we derive an inequality which can be exploited to verify the instability manifestation of non-supersymmetric four dimensional black holes via a characteristic function. For several species, we show that this function matches with the weak gravity swampland conjecture. Then, we deal with the cutoff issue as an interval estimation problem by putting a lower bound on the black hole mass scale matching with certain results reported in the literature. Using the developed formalism for the proposed instability scenarios, we present a suppression mechanism to the remnant production rate. Furthermore, we reconsider the stability study of the Reissner–Nordström black holes. As a result, we show that the approached instabilities prohibit naked singularity behaviors.
Keywords: Black holes, Swampland conjectures, Fermi ball, Hawking evaporation, Species scale.
1 Introduction
Recently, a special interest has been devoted to the study of effective field theories (EFTs) from higher dimensional supergravity scenarios such as superstrings and M-theory inspired models in connection with the swampland program [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. The latter aims to address such EFTs through discussions on the superstring landscape. Moreover, it could be used to check for consistencies with quantum gravity (QG). Moreover, this program could provide an alternative falsification mechanism in the light of the absence of the empirical evidence for certain physical theories. A close inspection shows that many swampland criteria have been suggested, developed, and refined in the context of superstring compactifications. Generally, these compactifications generate various scalar fields in the associated spectrum. Some of them correspond to appropriate geometric concepts including the size and the shape of internal compact spaces with non-trivial holonomy groups as well as the Calabi-Yau manifolds [11]. These scenarios have been largely elaborated in order to check and verify the application of the suggested swampland conditions. In contact with such developments, two relevant conjectures have been investigated being known by distance conjecture (DC) and weak gravity conjecture (WGC). The first one concerns the implication of the stringy moduli space in the EFT building models. In superstring theory, for instance, this space is parametrized by massless scalar fields originated from the compactification mechanism dealing with the dimension reduction of the space-time in which the supergravity spectrums live. This physical road could predict certain light particles and towers at large moduli space distances according to T-duality arguments and stringy size behaviors [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. The second conjecture (WGC) was first proposed in [25]. After that, it has been developed and approached from different angles[26, 27, 28, 29, 30, 31, 32]. This conjecture stems from the fact that gravity is the weakest fundamental interaction which is translated into two main observations. First, a cutoff scale at which the low energy EFT breaks is bound from the top by a gauge coupling. Secondly, the stable black hole horizons must have little to no charge compared to black ole mass. Otherwise at extremal limits the horizon become unstable, effectively suppressing naked singularities. Moreover, the correspondence between the WGC and the DC had been explained in [33, 34, 35, 36]. Specifically, the saturation of WGC bounds by extending the moduli distance is a common observation.
In black hole physics in arbitrary dimensions, WGC has been manipulated to put certain constraints on the charge per unit mass [35, 36, 37]. It has been also exploited to investigate small black holes in four dimensions derived from string theory using brane physics [38]. Moreover, the primordial black holes and the black hole species have been also approached in connection with WGC constraints. In particular, it has been revealed that the primordial extremal black holes with appropriate mass conditions could be considered as dark matter candidates [39, 40]. Beside that, the black holes evaporating down to the Planck scales have been studied in the light of the UV cutoff problems and the no-global symmetry conjecture [25].
The aim of this work is to contribute to these activities by approaching certain black hole issues, including remnants, by providing a statistical description relaying on WGC in the swampland program. Inspired by the Pauli exclusion principal in the context of the Fermi sphere, we derive an inequality which can be utilized to verify the instability manifestation of the black holes by help of a statistical characteristic function. For a generic number of species, we show that this function is in agreement with the weak gravity swampland conjecture. Next, we examine the possibility of using the cutoff as an interval estimation problem by imposing an upper limit on the black hole mass scale, which is compatible with some results obtained in the literature. [41, 42, 43, 44, 45, 46, 47, 48]. Using the developed formalism for the proposed instability scenarios, we provide a remnant production rate suppression mechanism. Moreover, we reconsider the stability behaviors of the Reissner–Nordström black holes. Precisely, we show that the proposed instabilities prohibit naked singularity behaviors.
The present work is structured as follows. In section 2, we present a concise discussion of black hole remnants. In section 3, we attempt to derive quantities allowing one to elaborate certain aspects of WGC and approach the remnant problems. In section 4, we present a swampland statistical description for the Reissner–Nordström black holes. The last section summarizes the main conclusion and provides some open questions.
2 Remnant production rate
In this section, we would like to address issues associated with the contemporary theory of black holes. In particular, we discuss the corresponding remnant problems. As the name hints, the remnants are evaporating black hole remains. They have been proposed to overcome the information loss problem in black hole physics. Various attempts have been made to solve this problem. Concretely, a black hole formation description was first proposed in [49] by establishing the expression form $\frac{M_{p}^{3}}{\pi^{3}}e^{-\frac{M_{p}^{2}}{16\pi T^{2}}}$, where $M_{p}$ is the Planck mass. Later, this work has been developed by providing a classical description of such a formation[50]. In these contributions, the formation rate of the black holes has been derived, at a finite temperature. According to the works developed in [51], these scenarios have been linked to the remnant production rate via the following form
$P\sim N_{R}e^{-\frac{M^{2}}{M_{p}^{2}}}$ (2.1)
where $N_{R}$ is the number of different remnant species. $M$ is the mass of the black hole. The main issue of this description is that the statistics of such objects do not add up. In fact, it has been observed that the production probability is very large for remnants at small mass scales including the Planck one. Alternatively, in order to not modify the Bekenstien-Hawking entropy, it has been argued that the remnant population could take the following form
$N_{R}\sim e^{2rM_{p}},$ (2.2)
where $r$ is the distance of a fiducial observer in the thermal atmosphere [52]. In this way, the remnants are taken to be in the Planck scale. However, other mass values could be considered. Additionally, at such small scales, the distance has been identified with the internal entropy of the black hole $S_{R}$ through the relation $r=\frac{S_{R}}{M_{p}}$. Therefore, the remnant production rate can be governed by the black hole internal entropy, being an increasing quantity, muddying the water further.
In what follows, we discuss some considerations of black holes, including remnants, by offering a statistical description based on the WGC within the swampland program.
3 WGC from the asymmetric instability
For letter use, we would like to present, first, a statistical scenario of the instability behaviors inspired by WGC being originally suggested to solve certain black hole problems including the remnant ones [25]. Roughly, this conjecture stands on two pillars, motivating the present study. First, it addresses EFT cutoffs $\Lambda$ by providing a QG obstruction to a vanishing value of the gauge coupling $g$, prohibiting the restoration of the abelian phase global symmetry [36]. This is ensured by implementing the following four dimensional inequality
$\Lambda\leq gM_{p}.$ (3.1)
Secondly, it could solve the naked singularity problems by predicting the decay of the extremal black hole into EFT stable particles. [25]. In gauge theories weakly coupled to gravity, there exist charged states of a mass $m$ and an electric charge $q$ satisfying the inequality
$\frac{q}{m}\geq\frac{Q}{M}|_{\mbox{extermal}}=\mathcal{O}(1)$ (3.2)
where $Q$ is the black hole charge [36]. Instead of continuing to evaporate and consequently becoming a naked singularity, this inequality ensures that the extremal black hole decays to preserve regular horizons described by the identity
$\frac{Q}{M}<\mathcal{O}(1).$ (3.3)
This WGC motivation could be exploited to address the remnant problems by introducing an instability analysis. This concerns the black hole Hawking evaporating down to the Planck scale.
In the present investigation, we attempt to approach the problem of remnants via WGC using a statistical description. To do so, we first reconsider the study of such a conjecture through a suggested asymmetric instability scenario. This discussion will be based on the following proposed statement. In a stable (non decaying) cavity, the number of the accessible charged states should not exceed the available number of massive states representing the entirety of the system. This statement has been inspired by the Pauli exclusion principal which has been exploited to introduce the asymmetric instability (or asymmetric energy) in the semi-empirical mass model [54], where the higher energy levels are occupied by less stable states. Motivated by the Fermi description of the energy of a ball of non relativistic $\frac{1}{2}$-spin fermions (Fermi ball) [55], we would like to provide a possible description of such instabilities. Concretely, this will be done using a Fermi spherical description. In this way, the number of charged states can be expressed as
$z=\iint\frac{\prod\limits^{2}_{\alpha=1}dx_{\alpha}dp_{\alpha}}{\hbar^{2}},$ (3.4)
where $z$ can be identified with the size parameter of the phase space described by the fibration $S^{2}\times\mathbb{R}^{2}$. It is also called the number density being a continuous parameter[55]. It is worth nothing that $S^{2}$ corresponds to the black hole surface and $\mathbb{R}^{2}$ is the momenta space of a particle in such a surface.
Using the expressions $dx_{1}dx_{2}=r^{2}\sin(\theta)d\theta d\phi=dA$ and $dp_{1}dp_{2}=dp_{\theta}dp_{\phi}=d^{2}p$ with the condition $p_{\theta}=p_{\phi}=p$, this gives arise the following Fermi sphere law
$\frac{dz}{dp}=\frac{Ap}{\hbar^{2}}.$ (3.5)
In the black hole physics, for instance, the quantity $A$ can be identified with the event horizon area of a Schwarzschild black hole, given by $A=\frac{16\pi G^{2}M^{2}}{c^{4}}$ [53]. Combining Eq.(3.5), and the relation $p=\sqrt{2m_{g}E_{z}}$, the energy associated with the asymmetric instability is found to be
$E=\int E_{z}dz=\frac{\hbar^{2}c^{4}z^{2}}{32\pi G^{2}M^{2}m_{g}},$ (3.6)
where $m_{g}$ is the mass of a unit charge (particle/specie) of a gauge coupling $g$. This quantity is needed to define the characteristic function being the inverse of the production rate of the specie $(g,m_{g})$ at a thermodynamic equilibrium. Indeed, it can be expressed as follows
$\rho=\frac{1}{\mathcal{Z}}e^{-\frac{\hbar^{2}c^{4}z^{2}}{32\pi k_{B}TG^{2}M^{2}m_{g}}}$ (3.7)
where $T$ is the temperature and $\mathcal{Z}$ is the corresponding partition function [54]. The normalization constraint of the characteristic function leads to
$\mathcal{Z}=\iint e^{-\frac{\hbar^{2}c^{4}z^{2}}{32\pi k_{B}TG^{2}M^{2}m_{g}}}\frac{r^{2}\sin(\theta)d\theta d\phi dp_{\theta}dp_{\phi}}{\hbar^{2}}.$ (3.8)
Using integration techniques, we obtain
$\mathcal{Z}=\frac{32\pi k_{B}TG^{2}M^{2}m_{g}}{z\hbar^{2}c^{4}}.$ (3.9)
The characteristic function given by Eq.(3.7) can be treated as a probability of accessing a state $(M,g,m_{g})$. In fact, the instability can be envisaged on the basis of decreasing values linked to the probability of accessing a flow of species produced in pairs on small scales. In order to examine such behaviors, we need to calculate the decay rate. Indeed, we get
$\frac{\partial\rho}{\partial(\frac{1}{M^{2}})}=\frac{zc^{4}\hbar^{2}}{36\pi k_{B}TG^{2}m_{g}}(1-\frac{c^{4}\hbar^{2}z^{2}}{36\pi k_{B}TG^{2}M^{2}m_{g}})e^{-\frac{\hbar^{2}c^{4}q^{2}}{32\pi k_{B}TG^{2}M^{2}m_{g}}}.$ (3.10)
The asymmetric instability corresponds to the constraint $\frac{\partial\rho}{\partial(\frac{1}{M^{2}})}<0$. This is insured by the inequality
$\frac{c^{4}\hbar^{2}z^{2}}{32\pi k_{B}TG^{2}M^{2}m_{g}}>1.$ (3.11)
Using the Planck units and the global charge, we have
$\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{M^{3}_{p}T_{p}}.$ (3.12)
Taking $M=NM_{p}$, the inequality corroborates the proposed statement up to a factor $\sqrt{\frac{32\pi m_{g}T}{M_{p}^{2}T_{p}}}$. Indeed, the proposed instability stems from the fact that the number of charged states $z$ exceeds the available number of massive states $N$. However, one should factor in the characteristics of the cavity, being the mass for the contained states $m_{g}$, the equilibrium temperature $T$, and the corresponding Planck scales.
Moreover, the instability of the states satisfying the inequality (3.12) increases when there are species maximizing the ratio $\frac{z}{m_{g}}$. This can be summarized by the following statement: A consistent EFT should always have species that maximize the ratio $\frac{|q_{i}|}{m_{i}}$, with $i=1,\ldots,N_{sp}$ where $N_{sp}$ is the number of the species and one has $q_{i}=gz_{i}$. Moreover, one considers that the black hole is composed of non-relativistic fermions with different integer spins. Thus, we can decompose the net charge $Q$ into elementary charges of species $N_{sp}$. According to [35], this has been formulated as follows
$\frac{|Q|}{M}\leq\frac{|\sum\limits_{i}q_{i}|}{\sum\limits_{i}m_{g}}\leq\frac{\sum\limits_{i}|q_{i}|}{\sum\limits_{i}m_{g}}\leq\max_{i}\frac{|q_{i}|}{m_{i}},$ (3.13)
describing a decay matching with the following conservation laws
$M=m_{g}+\Delta E,\hskip 6.0ptQ=\sum_{i}q_{i}$ (3.14)
where one has used $m_{g}=\sum\limits_{i}^{N_{sp}}m_{i}$. In the present study, furthermore, we assume that the particles of mass $m_{g}$ have an upper limit of species of maximal sizes with maximum possible discrete charges. A similar description has been elaborated for black holes, being endowed with large numbers of species associated with certain discreet symmetries [41]. However, it has been observed that it adapts better to a particle of mass $m_{g}$, by considering the inequality $\sum\limits_{i}m_{i}\leq\max_{i}m_{i}=N_{sp}M_{p}$. Handling Eq.(3.14) and (3.13) in the proposed characteristic function Eq.(3.7), we obtain the following inequality at some fixed black hole mass $M$
$\mbox{min }\rho\leq\frac{zM^{3}_{p}T_{p}}{32\pi M^{2}m_{g}T}e^{-\frac{M^{3}_{p}T_{p}}{32\pi\sum\limits_{i}m_{i}T}\frac{(\sum\limits_{i}z_{i})^{2}}{(\sum\limits_{i}m_{i})^{2}}}\leq\frac{zM^{3}_{p}T_{p}}{32\pi M^{2}m_{g}T}e^{-\frac{M^{3}_{p}T_{p}}{32\pi M^{2}T}\frac{(\sum\limits_{i}z_{i})^{2}}{\sum\limits_{i}m_{i}}},$ (3.15)
where one has used a minimal value of the characteristic function
$\mbox{min }\rho=\frac{zM^{3}_{p}T_{p}}{32\pi M^{2}m_{g}T}\exp\left(-\frac{M^{3}_{p}T_{p}}{32\pi\sum\limits_{i}m_{i}T}\left({\max_{i}}\left(\frac{z_{i}}{m_{i}}\right)\right)^{2}\right).$ (3.16)
In this way, the fastest decay channel is described by species maximizing the ratio $z_{i}/m_{i}$, in accordance with WGC. Exploiting, furthermore, the bound on $\sum\limits_{i}m_{i}$, we get an upper bound on the minimal value of the characteristic function
$\mbox{min }\rho\leq\frac{zM^{3}_{p}T_{p}}{32\pi M^{2}m_{g}T}\exp\left(-\frac{M^{2}_{p}T_{p}}{32\pi N_{sp}T}\left({\max_{i}}\left(\frac{z_{i}}{m_{i}}\right)\right)^{2}\right).$ (3.17)
In this bound on the minimal value, a cutoff term $\frac{M_{p}^{2}}{N_{sp}}$ has appeared. This allows one to confront the proposed instability behaviors with the previous works [41, 42, 43, 44, 45, 46, 47, 48]. The implications of such cutoffs will be elaborated in the forthcoming discussions.
3.1 UV divergence and cutoffs
In this part, we discuss some of the ambiguities surrounding the above cutoff points, in conjunction with WGC. To do so, we first reveal the UV divergence behavior in the proposed characteristic function. This divergence is encountered by summing over all the possible states in the phase space $\{A,p_{\phi},p_{\theta}\}$. This sum is an integral determined by the cumulative distribution function on the set of random variables $(A,p_{\phi},p_{\theta})$. The scenario, in which the black hole evaporates completely, is a possible interpretation to the convergence of such a function. In this context, the cumulative distribution function can be expressed as follows
$\mathbb{P}=\iint\rho\frac{dAd^{2}p}{\hbar^{2}}=\iint\frac{z}{12k_{B}Tm_{g}A}e^{-\frac{p^{2}}{2m_{g}k_{B}T}}dAd^{2}p.$ (3.18)
This integral reduces to
$\mathbb{P}=\int^{r_{h}}_{r_{0}}\frac{dr}{r}$ (3.19)
where $r_{h}$ is the even horizon radius of the involved black hole. It turns out that this integral diverges by sending $r_{0}$ to zero. To overcome such an issue, the sum needs a cut-off when the probability of accessing states with $r_{0}=0$ is zero. This argument could be motivated by the fact that the issue stems from the infinite sum over inaccessible states. To explore this issue further, we can treat the cutoff as an interval estimation problem. This could be done by considering the characteristic function as a non normalized gamma distribution given by
$\rho=X^{\alpha-1}e^{-\beta X},$ (3.20)
where $\alpha$ and $\beta$ are the shape and the scale parameters, respectively. $X$ is the random variable. In the present investigation $X$ is identified with $\frac{z\tau}{M^{2}}$ where one has used $\tau=\frac{M_{p}^{3}T_{p}}{m_{g}T}$. In order to avoid small amplitudes, we introduce the normalizing factor $\frac{\beta^{\alpha}}{\Gamma(\alpha)}$, to be useful. Taking $\alpha=2$ and $\beta=z$, we can calculate the variance of such a distribution. Precisely, we find
$Var[\frac{z\tau}{M^{2}}]=\frac{2}{z^{2}}.$ (3.21)
In this way, the mass can be expressed in terms of the standard deviation $\sqrt{Var[\frac{z\tau}{M^{2}}]}$ via the mean $\langle\frac{z\tau}{M^{2}}\rangle$ as follows
$M$ $=$ $\frac{\sqrt{z\tau}}{\sqrt{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}}=\frac{z\sqrt{\tau}}{\sqrt{2+n\sqrt{2}}}$ (3.22)
where $n$ is an integer describing the deviation degree. Accordingly, the amplitude of accessing a mass value in the rang $0,\frac{z\tau}{M^{2}}]$ is $\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$ 111 via the computation $\frac{z^{2}}{\Gamma(2)}\int_{0}^{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}Xe^{-zX}dX=\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$. where $\gamma(\alpha,x)$ is the lower incomplete gamma function given by the integral $\int^{x}_{0}x^{\alpha-1}e^{-x}dx$ [56]. Taking $n=20$, the amplitude reaches the value of $99.999\hskip 2.0pt999\hskip 2.0pt9997797\%$ for $M>\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$.Translating this for normal distributions, we conclude that the states bellow a cutoff $M_{0}=\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$ are inaccessible, with a $7\sigma$ confidence.
In the case of many species, Eq.(3.14) and Eq.(3.13) describe a mass scale cutoff at $\frac{\sum\limits_{i}z_{i}}{\sqrt{2+n\sqrt{2}}}\sqrt{\frac{M_{p}^{3}T_{p}}{32\pi\sum\limits_{i}m_{i}T}}$. Moreover, considering such a upper bound on the mass $\sum\limits_{i}m_{i}$ the defined cutoff in the proposed scenario should satisfy
$\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T(2+n\sqrt{2})}}(\sum\limits_{i}z_{i})<M_{0}.$ (3.23)
One can choose $N_{sp}$ to be large enough so that the term $2+n\sqrt{2}$ is ignored. According to [41], this implies that one has $\Lambda_{sp}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $\Lambda_{sp}$ represents the species scale. This observation is consistent with the inverse relation between the species scale $\Lambda_{sp}$ and the number of species $N_{sp}$, which could be linked to the gauge coupling via $N_{sp}\sim\frac{1}{g^{2}}$, restoring the condition given in Eq.(3.1). However, a correction term $\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})$ has appeared.
Now, we check the consistency of WGC with the assumptions that the black holes with weak gauge couplings are the most common. To do so, we compute the probability of accessing states with either $M>\frac{Q\sqrt{\tau}}{g}$ or $M<\frac{Q\sqrt{\tau}}{g}$, defined by $\mathcal{P}=\frac{1}{\langle M\rangle}\int^{a}_{b}\frac{z\tau}{M^{2}}e^{-\frac{z\tau}{M^{2}}}dM$. In order to get such probabilities, we should integrate along the interval $\frac{Q\sqrt{\tau}}{g},\infty[$. Taking the black hole mass as the measure, the probability of accessing states with $M>\frac{Q\sqrt{\tau}}{g}$ is found to be
$\mathcal{P}(M>\frac{Q\sqrt{\tau}}{g})=erf(1).$ (3.24)
However, the probability of accessing states with $Q\sqrt{\tau}>M$ is
$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=cerf(1)$ (3.25)
where one has used $erf(1)=\frac{2}{\sqrt{\pi}}\int^{1}_{0}e^{-x^{2}}dx$ and cerf(x)=1erf(x)cerf(x)=1-erf(x)[56]. It follows from Eq.(3.24) and Eq.(3.25) that one has
$\mathcal{P}(M>\frac{Q\sqrt{\tau}}{g})>\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g}).$ (3.26)
This shows that WGC is consistent with the aforementioned assumptions. In connection with the previous arguments about cutoffs, the probability $\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})$ decreases by sending $M_{0}$ to $\frac{Q\sqrt{\tau}}{g}$. This can be supported by the relation
$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=Erf(\frac{Q\sqrt{\tau}}{gM_{0}})-Erf(1),$ (3.27)
matching with Eq.(3.23).
3.2 Remnant production rate and WGC
At this stage, we address the implication of the developed characteristic function in the remnant production rate. In black hole physics, the probability of entering a $(N_{R},M,g,m_{g})$ state is determined as a function of the rate of remnant production and the characteristic function associated with the asymmetric instability. Explicitly, such a probability is found to be
$P(P_{BH}\cup\rho)=\frac{zM^{3}_{p}T_{p}}{32\pi M^{2}m_{g}T}e^{2\pi rM_{p}-\frac{M^{2}}{M_{p}^{2}}-\frac{M^{3}_{p}T_{p}z^{2}}{32\pi M^{2}m_{g}T}}.$ (3.28)
At Planck scales, the probability reduces to the following form
$P(P_{BH}\cup\rho)\sim ze^{-z^{2}}.$ (3.29)
Thus, it is clear that the asymmetric instability suppresses the remnant production rate.
The suppression of accessible states could be interpreted for small scale regimes. As the black hole evaporates down to a small scale, the horizon size decreases. Due to the Pauli exclusion principal, the pair produced charges can no longer exist on the black hole surface. However, the existence of such objects can be supported by a disintegrating scenario to other stable black holes and particles matching with EFTs in question. This should follow possible decay paths with different probabilities. For a black hole evaporating down to a Planck scale, the only disintegration path is to decay into stable particles as stated by WGC [25, 36].
4 Swampland statistics for charged black holes
As applications, the present investigation automatically motivates the discussion of a charged solution known by the Reissner–Nordström black hole in four dimensions [53]. This black hole is described by the following metric line
$ds^{2}=f(r)dt^{2}-\frac{1}{f(r)}dr^{2}-r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (4.1)
For this solution, one has considered
$f(r)=1-\frac{2GM}{c^{2}r}+\frac{Q^{2}G}{4\pi\epsilon_{0}c^{4}r^{2}}$ (4.2)
where $\epsilon_{0}$ is the vacuum electric permeability. According to [53], the event horizon can be obtained by solving the constraint $f(r_{h})=0$. Indeed, two solutions can be derived
$r_{h}^{\pm}=\frac{G}{c^{2}}\left(M\pm\sqrt{M^{2}-\frac{Q^{2}}{G\pi\epsilon_{0}}}\right).$ (4.3)
According to previous discussions, the black hole could involve, either a net charge over time due to the Hawking radiation (a charge from its formation) or both scenarios. For a generic situation, the asymmetric probability densities are found to be
$\rho^{+}$ $=$ $\frac{z\tau}{M^{2}(1+\sqrt{1-\frac{Q^{2}}{G\pi\epsilon_{0}M^{2}}})^{2}}e^{-\frac{\tau z^{2}}{M^{2}(1+\sqrt{1-\frac{Q^{2}}{G\pi\epsilon_{0}M^{2}}})^{2}}}$ (4.4)
$\rho^{-}$ $=$ $\frac{z\tau}{M^{2}(1-\sqrt{1-\frac{Q^{2}}{G\pi\epsilon_{0}M^{2}}})^{2}}e^{-\frac{\tau z^{2}}{M^{2}(1-\sqrt{1-\frac{Q^{2}}{G\pi\epsilon_{0}M^{2}}})^{2}}}.$ (4.5)
In our case, we only focus on $\rho^{+}$, while a similar analysis is possible for $\rho^{-}$. Concretely, the inequality (3.12) becomes
$\frac{Q}{M}>\frac{g}{\sqrt{{\tau}}}\left(1+\sqrt{1-\frac{Q^{2}}{G\pi\epsilon_{0}M^{2}}}\right).$ (4.6)
For real behaviors of the inequality, one should have $\frac{Q}{M}<G\pi\epsilon_{0}$. This is insured by the inequality
$\frac{g^{2}}{\tau}<G\pi\epsilon_{0}.$ (4.7)
For more clarifications the inequality (4.6) can be simplified to
$\frac{Q}{M}>\lambda=\frac{2\frac{g}{\sqrt{\tau}}}{1+\frac{g^{2}}{\tau G\pi\epsilon_{0}}}.$ (4.8)
The condition $\sqrt{G\pi\epsilon_{0}}>\lambda$ is thus insured by the in inequality (4.7). It is worth noting that in terms of charged particle mass, at $T=T_{p}$, the inequality (4.7) generates the upper bound
$m_{g}<\frac{M_{p}}{168\pi\alpha}$ (4.9)
with $\alpha$ the fine structure constant. This bound has been observed to be adhered to
by the standard model. Therefore, the following Table.(1) contain the different black hole horizon scenarios.
Black hole Horizon Stable Unstable
Regular horizon $\sqrt{G\pi\epsilon_{0}}>\lambda>\frac{Q}{M}$ $\sqrt{G\pi\epsilon_{0}}>\frac{Q}{M}>\lambda$
Extremal horizon Contradicts inequality (4.7) $\sqrt{G\pi\epsilon_{0}}=\frac{Q}{M}>\lambda$
Naked singularity Contradicts inequality (4.7) $\frac{Q}{M}>\sqrt{G\pi\epsilon_{0}}>\lambda$
Table 1: The main stability scenarios for different black hole horizons
This means that the black hole should disintegrate due to asymmetric instability before reaching its extreme limit, for which no stable solution is possible at this level, as shown in Table (1). Thus, the naked singularity scenarios can be avoided in accordance with the WGC arguments [25]. Indeed, we discern between two limits. One corresponds to a situation where there is no room to fit more charged states. The second one asserts that the black hole solution exhibits a naked singularity behavior. Both situations are governed by a different charge to the mass ratio inequalities. These assertions would be easily verified in the case of extremal black holes. Evidently, in order to be unstable, such black holes should satisfy
$\frac{Q}{M}>\frac{g}{\sqrt{{\tau}}}.$ (4.10)
Considering the fact that $\frac{Q}{M}=\sqrt{G\pi\epsilon_{0}}>\frac{g}{\sqrt{{\tau}}}$, we obtain a consistent result with Eq.(4.7).
Regarding the remnant production rate, in this scenario, we assume that the black hole evaporates down to a Planck scale by keeping a regular horizon. Considering the Reissner–Nordström black holes at such small scales, we observe that the probability of accessing a state $(N_{R},M,Q,g,m_{g})$ can take the following form
$P(P_{BH}\cup\rho_{+})\sim\rho_{+}.$ (4.11)
Consequently, the asymmetric instabilities in these black hole solutions allow one to reach the desired result being the suppression of the remnant production rate.
5 Conclusion
In this paper, we have approached certain black hole issues including remnant problems by providing a statistical description based on WGC in the swampland program. It has been suggested that this conjecture basically postulates that, at small scales, the black holes should be unstable. In fact, there are possible decay channels induced by set instabilities. Based on such arguments, we have carried out a statistical investigation for non-supersymmetric four dimensional black holes, where such instabilities may be underlined by the Pauli exclusion principal. Precisely, we have shown that this statistical framework allows one to derive the charge to the mass ratio inequality $\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{T_{p}M^{3}_{p}}$, where $m_{g}$ is the mass of a particle with a gauge coupling $g$ and $T$ is a thermodynamical equilibrium temperature with Planck scales. Consequently, it has been remarked that this formulation is similar to that of the weak gravity swampland conjecture, as the proposed instabilities also appear at small scales. Supported by the proposed instabilities, we have provided a suppression mechanism to the remnant production rate. Moreover, we have considered a generic situation to address the species scale. Specifically, we have treated the cuts in the mass scale of the black hole $M_{0}$ as an interval estimation problem. For a generic situation, this has resulted in a black hole mass cutoff lower bound $\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $z_{i}$ and $N_{sp}$ are the number of charged states and the number species, respectively. We have observed that the upper bound can be considered as a corrected form of the species scale introduced in the literature. Using the developed formalism of the proposed instabilities, we have reconsidered the study of the Reissner–Nordström black holes as a possible application for charged solutions. As a result, we have found that this illustrating model is consistent with the previous observations, including the remnant production rate suppression. Furthermore, we have revealed that the proposed instabilities prohibit naked singularity behaviors due to the constraint $\frac{g^{2}}{\tau}\leq G\pi\epsilon_{0}$.
This work leaves certain open issues. A natural question concerns extra black hole backgrounds including either the rotating parameter or extra scalar fields originated from different scenarios. A possible road is to implement stringy scalars which could be derived from superstring compactification mechanisms via internal compact spaces with non-trivial holonomy groups such as Calabi-Yau geometries.

### v3

@@
 As a result, we show that the approached instabilities prohibit naked singularity behaviors.
-Keywords: Black holes, Swampland conjectures, Fermi ball, Hawking evaporation, Species scale.
+Keywords: Black holes, Swampland conjectures, Fermi sphere, Hawking evaporation, Species scale.
 1 Introduction
@@
 Generally, these compactifications generate various scalar fields in the associated spectrum.
-Some of them correspond to appropriate geometric concepts including the size and the shape of internal compact spaces with non-trivial holonomy groups as well as the Calabi-Yau manifolds [11].
+Some of them correspond to appropriate geometric concepts including the size and the shape of internal compact spaces with non-trivial holonomy groups as well as the Calabi-Yau manifolds [11, 12, 13, 14].
 These scenarios have been largely elaborated in order to check and verify the application of the suggested swampland conditions.
@@
 The aim of this work is to contribute to these activities by approaching certain black hole issues, including remnants, by providing a statistical description relaying on WGC in the swampland program.
-Inspired by the Pauli exclusion principal in the context of the Fermi sphere, we derive an inequality which can be utilized to verify the instability manifestation of the black holes by help of a statistical characteristic function.
+Inspired by the Pauli exclusion principle in the context of the Fermi sphere, we derive an inequality which can be utilized to verify the instability manifestation of the black holes by help of a statistical characteristic function.
 For a generic number of species, we show that this function is in agreement with the weak gravity swampland conjecture.
@@
 Later, this work has been developed by providing a classical description of such a formation[50].
-In these contributions, the formation rate of the black holes has been derived, at a finite temperature.
+In these contributions, the formation rate of the black holes has been derived, at a finite temperature being related to the black hole mass via the Hawking radiation, $T=\frac{M_{p}^{2}}{8\pi M}$.
 According to the works developed in [51], these scenarios have been linked to the remnant production rate via the following form
@@
 In fact, it has been observed that the production probability is very large for remnants at small mass scales including the Planck one.
-Alternatively, in order to not modify the Bekenstien-Hawking entropy, it has been argued that the remnant population could take the following form
+In any case, the number of remaining species can be infinite, as there is no fixed limit for this number.
+In certain cases, it has been observed that the situation can be exacerbated when the internal entropy of the remains is used to count $N_{R}$.
+In order to avoid altering the Bekenstien-Hawking entropy, it has been proposed that the remaining population could take the following form
 $N_{R}\sim e^{2rM_{p}},$ (2.2)
@@
 In this way, the remnants are taken to be in the Planck scale.
-However, other mass values could be considered.
+However, lower mass values could be considered.
 Additionally, at such small scales, the distance has been identified with the internal entropy of the black hole $S_{R}$ through the relation $r=\frac{S_{R}}{M_{p}}$.
 Therefore, the remnant production rate can be governed by the black hole internal entropy, being an increasing quantity, muddying the water further.
+From another perspective, in [53], it was pointed out that when the content of a black hole is separated into several quantum states, called “informons,” the decay rate of a weakly coupled electric field into these informons (the rate of residue production) includes a divergent term in the form of a sum of exponential expressions over the number of these states.
+This divergence would be caused by the infinite number of informons whose mass is less than the Planck scale.
+As a prelude to what follows, this question will be addressed by proposing that it is not possible to fit an infinite number of states into the black hole.
+Alternatively, the usual Hawking evaporation would be interrupted by the decay of the black hole along certain probable decay channels.
 In what follows, we discuss some considerations of black holes, including remnants, by offering a statistical description based on the WGC within the swampland program.
@@
 [25].
-In gauge theories weakly coupled to gravity, there exist charged states of a mass $m$ and an electric charge $q$ satisfying the inequality
+Explicitly, in gauge theories weakly coupled to gravity, there exist charged states of a mass $m$ and an electric charge $q$ satisfying the inequality
 $\frac{q}{m}\geq\frac{Q}{M}|_{\mbox{extermal}}=\mathcal{O}(1)$ (3.2)
@@
 In a stable (non decaying) cavity, the number of the accessible charged states should not exceed the available number of massive states representing the entirety of the system.
-This statement has been inspired by the Pauli exclusion principal which has been exploited to introduce the asymmetric instability (or asymmetric energy) in the semi-empirical mass model [54], where the higher energy levels are occupied by less stable states.
-Motivated by the Fermi description of the energy of a ball of non relativistic $\frac{1}{2}$-spin fermions (Fermi ball) [55], we would like to provide a possible description of such instabilities.
+This statement has been inspired by the Pauli exclusion principle which has been exploited to introduce the asymmetric instability (or asymmetric energy) in the semi-empirical mass model [54], where the higher energy levels are occupied by less stable states.
+Motivated by the Fermi description of the energy of a ball of non relativistic $\frac{1}{2}$-spin fermions [55], we would like to provide a possible description of such instabilities.
 Concretely, this will be done using a Fermi spherical description.
@@
 It is worth nothing that $S^{2}$ corresponds to the black hole surface and $\mathbb{R}^{2}$ is the momenta space of a particle in such a surface.
-Using the expressions $dx_{1}dx_{2}=r^{2}\sin(\theta)d\theta d\phi=dA$ and $dp_{1}dp_{2}=dp_{\theta}dp_{\phi}=d^{2}p$ with the condition $p_{\theta}=p_{\phi}=p$, this gives arise the following Fermi sphere law
+In the following, we use the expressions $dx_{1}dx_{2}=r^{2}\sin(\theta)d\theta d\phi=dA$ and $dp_{1}dp_{2}=dp_{\theta}dp_{\phi}=d^{2}p$ with the condition $p_{\theta}=p_{\phi}=p$.
+By deriving the number of states with respect to the momentum and integrating over the spatial variables, we arrive at the following Fermi sphere law
 $\frac{dz}{dp}=\frac{Ap}{\hbar^{2}}.$ (3.5)
+This law basically dictate how one should count the states in the given phase space.
 In the black hole physics, for instance, the quantity $A$ can be identified with the event horizon area of a Schwarzschild black hole, given by $A=\frac{16\pi G^{2}M^{2}}{c^{4}}$ [56].
+Since we wish to establish a scenario of instability governed by a charge-to-mass ratio, we will not delve into more detailed phenomenological studies of these states.
+We will therefore consider that the observer counting the states is in a reference frame where relativistic effects are negligible.
 Combining Eq.(3.5), and the relation $p=\sqrt{2m_{g}E_{z}}$, the energy associated with the asymmetric instability is found to be
@@
 $\frac{c^{4}\hbar^{2}z^{2}}{32\pi k_{B}TG^{2}M^{2}m_{g}}>1.$ (3.11)
-Using the Planck units and the global charge, we have
+Considering quantized total charge $Q=zg$, and using the Planck units and the global charge, we have
 $\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{M^{3}_{p}T_{p}}.$ (3.12)
@@
 $Var[\frac{z\tau}{M^{2}}]=\frac{2}{z^{2}}.$ (3.21)
-In this way, the mass can be expressed in terms of the standard deviation $\sqrt{Var[\frac{z\tau}{M^{2}}]}$ via the mean $\langle\frac{z\tau}{M^{2}}\rangle$ as follows
-$M$ $=$ $\frac{\sqrt{z\tau}}{\sqrt{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}}=\frac{z\sqrt{\tau}}{\sqrt{2+n\sqrt{2}}}$ (3.22)
-where $n$ is an integer describing the deviation degree.
-Accordingly, the amplitude of accessing a mass value in the rang $0,\frac{z\tau}{M^{2}}]$ is $\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$ 111 via the computation $\frac{z^{2}}{\Gamma(2)}\int_{0}^{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}Xe^{-zX}dX=\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$.
+In this way, the mass can be treated as a step indexed by an integer.
+Thus it is expressed in terms of the standard deviation $\sqrt{Var[\frac{z\tau}{M^{2}}]}$ via the mean $\langle\frac{z\tau}{M^{2}}\rangle$ as follows
+$M_{n}$ $=$ $\frac{\sqrt{z\tau}}{\sqrt{\langle\frac{z\tau}{M^{2}}\rangle+n\sqrt{Var[\frac{z\tau}{M^{2}}]}}}=\frac{z\sqrt{\tau}}{\sqrt{2+n\sqrt{2}}}$ (3.22)
+where $n$ is an integer indicating the degree of deviation from the mean.
+Accordingly, the amplitude of accessing a mass value in the rang $0,\frac{z\tau}{M^{2}}]$ is $\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$ 111via the computation $\frac{z^{2}}{\Gamma(2)}\int_{0}^{\frac{2+n\sqrt{2}}{z}}Xe^{-zX}dX=\frac{\gamma(2,2+n\sqrt{2})}{\Gamma(2)}$.
 where $\gamma(\alpha,x)$ is the lower incomplete gamma function given by the integral $\int^{x}_{0}x^{\alpha-1}e^{-x}dx$ [57].
-Taking $n=20$, the amplitude reaches the value of $99.999\hskip 2.0pt999\hskip 2.0pt9997797\%$ for $M>\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$.Translating this for normal distributions, we conclude that the states bellow a cutoff $M_{0}=\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$ are inaccessible, with a $7\sigma$ confidence.
+As a simple illustration, Fig.(1) shows the selected sample by considering a degree of deviation $n=1$.
+Precisely, with such a degree of deviations, one can access a state in the range $\frac{M_{1}}{\sqrt{\tau}},\infty[$ with a probability of $85.5\%$.
+Figure 1: Cutoff mas $M_{1}$ at a degree $n=1$ in the $z=2$ case.
+Additionally, by taking $n=20$, the amplitude reaches the value of $99.999\hskip 2.0pt999\hskip 2.0pt9997797\%$ for $M\geq\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$.Translating this for normal distributions, we conclude that the states bellow a cutoff $M_{20}=\frac{z\sqrt{\tau}}{\sqrt{2+20\sqrt{2}}}$ are inaccessible, with a $7\sigma$ confidence.
 In the case of many species, Eq.(3.14) and Eq.(3.13) describe a mass scale cutoff at $\frac{\sum\limits_{i}z_{i}}{\sqrt{2+n\sqrt{2}}}\sqrt{\frac{M_{p}^{3}T_{p}}{32\pi\sum\limits_{i}m_{i}T}}$.
 Moreover, considering such a upper bound on the mass $\sum\limits_{i}m_{i}$ the defined cutoff in the proposed scenario should satisfy
-$\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T(2+n\sqrt{2})}}(\sum\limits_{i}z_{i})<M_{0}.$ (3.23)
+$\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T(2+n\sqrt{2})}}(\sum\limits_{i}z_{i})<M_{n}.$ (3.23)
 One can choose $N_{sp}$ to be large enough so that the term $2+n\sqrt{2}$ is ignored.
-According to [41], this implies that one has $\Lambda_{sp}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $\Lambda_{sp}$ represents the species scale.
+According to [41], this implies that one has $\Lambda_{sp}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{n}$, where $\Lambda_{sp}$ represents the species scale.
 This observation is consistent with the inverse relation between the species scale $\Lambda_{sp}$ and the number of species $N_{sp}$, which could be linked to the gauge coupling via $N_{sp}\sim\frac{1}{g^{2}}$, restoring the condition given in Eq.(3.1).
@@
 This shows that WGC is consistent with the aforementioned assumptions.
-In connection with the previous arguments about cutoffs, the probability $\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})$ decreases by sending $M_{0}$ to $\frac{Q\sqrt{\tau}}{g}$.
+In connection with the previous arguments about cutoffs, the probability $\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})$ decreases by sending a cutoff $M^{\star}$ to $\frac{Q\sqrt{\tau}}{g}$.
 This can be supported by the relation
-$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=Erf(\frac{Q\sqrt{\tau}}{gM_{0}})-Erf(1),$ (3.27)
+$\mathcal{P}(M<\frac{Q\sqrt{\tau}}{g})=erf(\frac{Q\sqrt{\tau}}{gM^{\star}})-erf(1),$ (3.27)
 matching with Eq.(3.23).
@@
 As the black hole evaporates down to a small scale, the horizon size decreases.
-Due to the Pauli exclusion principal, the pair produced charges can no longer exist on the black hole surface.
-However, the existence of such objects can be supported by a disintegrating scenario to other stable black holes and particles matching with EFTs in question.
+Due to the Pauli exclusion principle, the pair produced charges can no longer exist on the black hole surface.
+However, the existence of such objects can be supported by a disintegrating scenario into other stable black holes and particles matching with EFTs in question.
 This should follow possible decay paths with different probabilities.
-For a black hole evaporating down to a Planck scale, the only disintegration path is to decay into stable particles as stated by WGC [25, 36].
+For a black hole evaporating down to a Planck scale, the only disintegration path is to decay into stable particles as stated by WGC [25, 36].Lastly, the density function introduced should contribute to the Bekenstein-Hawking entropy, as shown in equation (3.28).
+This leads to significant results concerning the Bekenstein bound, where states are counted over a cubical box.
+However, the implications of this work for entropy will be the subject of future investigations.
 4 Swampland statistics for charged black holes
@@
 In fact, there are possible decay channels induced by set instabilities.
-Based on such arguments, we have carried out a statistical investigation for non-supersymmetric four dimensional black holes, where such instabilities may be underlined by the Pauli exclusion principal.
+Based on such arguments, we have carried out a statistical investigation for non-supersymmetric four dimensional black holes, where such instabilities may be underlined by the Pauli exclusion principle.
 Precisely, we have shown that this statistical framework allows one to derive the charge to the mass ratio inequality $\frac{Q^{2}}{M^{2}}>\frac{32\pi g^{2}m_{g}T}{T_{p}M^{3}_{p}}$, where $m_{g}$ is the mass of a particle with a gauge coupling $g$ and $T$ is a thermodynamical equilibrium temperature with Planck scales.
@@
 Specifically, we have treated the cuts in the mass scale of the black hole $M_{0}$ as an interval estimation problem.
-For a generic situation, this has resulted in a black hole mass cutoff lower bound $\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{0}$, where $z_{i}$ and $N_{sp}$ are the number of charged states and the number species, respectively.
+For a generic situation, this has resulted in a black hole mass cutoff lower bound $\frac{M_{p}}{\sqrt{N_{sp}}}\sqrt{\frac{T_{p}}{32\pi T}}(\sum\limits_{i}z_{i})<M_{n}$, where $z_{i}$ and $N_{sp}$ are the number of charged states and the number species, respectively.
 We have observed that the upper bound can be considered as a corrected form of the species scale introduced in the literature.
@@
 As a result, we have found that this illustrating model is consistent with the previous observations, including the remnant production rate suppression.
-Furthermore, we have revealed that the proposed instabilities prohibit naked singularity behaviors due to the constraint $\frac{g^{2}}{\tau}\leq G\pi\epsilon_{0}$.
+Furthermore, we have revealed that the proposed instabilities prohibit naked singularity behaviors due to the constraint $\frac{g^{2}}{\tau}<G\pi\epsilon_{0}$.
 This work leaves certain open issues.
@@
 A possible road is to implement stringy scalars which could be derived from superstring compactification mechanisms via internal compact spaces with non-trivial holonomy groups such as Calabi-Yau geometries.
+Conflicts of interest
+The authors declare that they have no conflicts of interest.